10. In the Keynesian Cross analysis, if the consumption function is given by C = 100 + 0.6(Y − T) and planned investment is 100, G = T = 100, then equilibrium Y is:

a) 350

b) 400

c) 600

d) 750

11. Assume that the money demand function is L(r) = 2200 − 200r, where r is the interest rate in percent. The money supply is 2000 and the price level is 2. The equilibrium interest rate is percent.

a) 2

b) 4

c) 6

d) 8

12. The IS curve shifts when all of the following economic variables change except:

a) the interest rate. b) government spending. c) tax rates. d) the marginal propensity to consume.
Answer #1

Q10
Answer

the equilbrium is at
Y=C+I+G
Y=100+0.6(Y-100)+100+100
Y=300+0.6Y-60
0.4Y=240
Y=600

Option c
=============
Q11
Answer
the equilibrium is at
M/P=L
2000/2=2200-200r
200r=2200-1000
200r=1200
r=6%

Option c
==========
Q12
Answer
Option a
the IS curve shifts if there is a change in all other option except the interest rate as the change in interest rate is a movement along the IS curve as the IS curve is a inverse realationship between interest rate and output.